write the primitive or anti derivative of f(x) = sqrt(x) + 1/sqrt(x)

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (x+\frac{1}{\sqrt{x}})dx
I = \int xdx+\int \frac{1}{\sqrt{x}}dx
I = \frac{x^2}{2}+\int \frac{1}{\sqrt{x}}dx
I = \frac{x^2}{2} + 2\sqrt{x} + c